OpenStudy (anonymous):

Write an equation for the line that passes through the given point with the given slope m. (10,1);m=1/5

OpenStudy (shamim):

Required formula is y-y1=m(x-x1)

OpenStudy (shamim):

x1,y1,m given

OpenStudy (.cheesecake.):

\[\huge y-y_1=m(x-x_1)\] where m=slope (x1,y1)